What is the resistance of a laptop cord that uses 120 volts and 1.6 amperes? Show all work including formulas and units.
Answer #1

Potential difference between the ends of the laptop cord, E = 120 V

Current through the cord, I = 1.6 A

So, resistance of the laptop cord, R = E/I

= 120 / 1.6 = 75.0 (Answer)